OpenStudy (anonymous):

help on finding the equation for a = 4, b = 4, c =. pleaz

OpenStudy (anonymous):

what happens

OpenStudy (anonymous):

c=4

OpenStudy (anonymous):

i said equation by the way pleaz

OpenStudy (eric_d):

Use pythagorean method

OpenStudy (eric_d):



OpenStudy (anonymous):

\(a^2 + b^2 = c^2\) \(4^2 + 4^2 = c^2\) What's 4 * 4?

OpenStudy (anonymous):

i just need equation i couldnt find it

OpenStudy (anonymous):

Oh, okay

OpenStudy (anonymous):

null last pleaz my eye be blind fer a sec there i do need to solve

OpenStudy (anonymous):

?

OpenStudy (anonymous):

means i didnt see where it said i had to solve pleaz help